Find the value of R:(9,r)(6,3)m=-1/3

Answer by ewatrrr(24785)

 
Hi,
(9,r)
(6,3) m= (r-3)/3 = -1/3
(r-3)/3 = -1/3
r-3 = -1
r+=+2
(9,2)
(6,3) m = -1/3